# On the Duality between Gradient Transformations and Adapters

### v1

@@ Abstract
We study memory-efficient optimization of neural networks with linear gradient transformations, where the gradients are linearly mapped to a lower dimensional space than the full parameter space, thus saving memory required for gradient accumulation and optimizer state persistence. The model parameters are updated by first performing an optimization step in the lower dimensional space and then going back into the original parameter space via the linear map’s transpose. We show that optimizing the model in this transformed space is equivalent to reparameterizing the original model through a linear adapter that additively modifies the model parameters, and then only optimizing the adapter’s parameters. When the transformation is Kronecker-factored, this establishes an equivalence between GaLore (Zhao et al., 2024) and one-sided LoRA (Hu et al., 2022). We show that this duality between gradient transformations and adapter-based reparameterizations unifies existing approaches to memory-efficient training and suggests new techniques for improving training efficiency and memory use.
@@ 1 Introduction
Training neural networks, in particular large language models (LLMs), can be extremely memory-intensive. Standard approaches for LLM training use gradient accumulation across multiple batches and optimizers such as Adam (Kingma & Ba, 2015), which maintains estimates of the first and second moments of the (stochastic) gradient. The amount of GPU memory needed for standard training is then roughly four times the amount of memory needed to store the model (assuming the gradients/optimization states are kept in the same precision as the model parameters).
@@
Our empirical experiments study this connection between linear gradient transformations and adapter-based reparameterizations in the context of memory-efficient LLM training. First, we perform a comparison across gradient projection-based and LoRA-based approaches for memory-efficient training and find that randomly sketching gradients works particularly well. We then exploit the adapter view of projected-gradient training [^1] Note that this notion of performing gradient descent with projections of the gradient is distinct from projected gradient descent (PGD) from the optimization literature. by developing a QLoRA-style (Dettmers et al., 2023) approach to GaLore-style training. We next show that the gradient projection view of LoRA adapters can improve distributed training of LLMs with parallel LoRA adapters (Huh et al., 2024) by suggesting an initialization scheme of worker-specific LoRA adapters tailored for distributed training. These results collectively demonstrate that this duality between linear gradient transformations and adapter-based reparameterizations is a productive lens with which to view neural network optimization, since it unifies several existing approaches and suggests new techniques for improving training efficiency and performance.
@@ 2.1 Memory Characteristics of LLM Training
Model parameters. We must keep the model’s parameters in memory, since these are used in various stages of the training process (e.g., to compute gradients). Here, it is useful to distinguish trainable parameters (which get updated regularly during training) from non-trainable parameters, which are not updated during training but may still be used in gradient computation.
@@
This work is mostly concerned with training LLMs in memory-constrained regimes where the model, optimizer, and gradient memory dominate, since activation storage can be made small by, e.g., reducing the (micro) batch size.
@@ 2.2 LoRA and GaLore
GaLore provides an alternative approach to memory-efficient pretraining. Instead of reparameterizing the weights to be a combination of full-rank and low-rank matrices—which increases the number of model parameters—GaLore performs a low-rank compression of the gradient matrix $\overline{{\mathbf{W}}}$ instead. The optimizer update is performed in this lower dimensional space, and the updated parameters are uncompressed back into the full parameter space before the next forward pass. Specifically, given a gradient matrix $\overline{{\mathbf{W}}}\in\mathbb{R}^{m\times n}$, GaLore uses a matrix ${\mathbf{P}}\in\mathbb{R}^{k\times m}$ (with $k<m$) to transform the gradient via ${\mathbf{P}}\overline{{\mathbf{W}}}\in\mathbb{R}^{k\times n}$, feeds this compressed gradient into a regular optimizer to obtain a pseudo-parameter update $\Delta\in\mathbb{R}^{k\times n}$, and then updates the original parameters via ${\mathbf{P}}^{\top}{\Delta}$. In practice, ${\mathbf{P}}$ is given by the top singular vectors of $\overline{{\mathbf{W}}}$, where in order to amortize the cost of SVD, ${\mathbf{P}}$ is updated only every so often. As with LoRA, GaLore reduces the memory needed to store the optimizer states, since optimization happens in the lower dimensional space.
@@ 3 Duality between Linear Gradient Transformations and Adapters
In this section, we prove that training a neural network using linear transformations of the gradient is equivalent to reparameterizing the neural network using specific linear adapters. We begin with the general case, where all parameters are treated as arbitrary vectors (§ 3.1). We then show how applying a Kronecker-factored linear transformation to the gradients of linear layers of the network is equivalent to training the model with a version of LoRA which inserts a trainable square matrix between the LoRA matrices (§ 3.2). From this, we further show that specializing to a specific choice of Kronecker-factored transformation establishes an equivalence between GaLore (Zhao et al., 2024) and “one-sided” LoRA (Hu et al., 2022) where one of the LoRA adapters is initialized in a particular way and remains frozen (§ 3.2); this recovers the equivalence established in recent work (Hao et al., 2024; Loeschcke et al., 2024).
@@ 3.1 General Case
Let $f({\mathbf{X}};\Theta)$ be a neural network over input ${\mathbf{X}}$ with trainable parameters $\Theta\in\mathbb{R}^{d}$, and further let $\overline{\Theta}\in\mathbb{R}^{d}$ be the gradient of some differentiable loss function ${\mathcal{L}}$ of the network $f$ with respect to $\Theta$, computed on a random data minibatch. We use the superscript $\left(\cdot\right)^{(t)}$ to specify a particular quantity’s value after $t$ optimizer steps, e.g., $\Theta^{(t)}$ are the network’s parameters after $t$ optimizer steps. We are also interested in the optimization trajectory of a model $(\Theta^{(0)},\Theta^{(1)},\ldots,\Theta^{(t)})$. Our results show that two optimization trajectories—one from training with gradient projections, and one from training with adapters—are equivalent.
@@
Because the above is adapting a neural network indirectly via another vector $\Lambda$ that is linearly mapped to the original parameter space, we refer to this as using a linear adapter, akin to the usage of “adapter” in the parameter-efficient finetuning literature (Houlsby et al., 2019; Hu et al., 2022). Since ${\mathbf{S}}^{\top}\in\mathbb{R}^{d\times r}$, we have $\dim(\Delta_{\Lambda}^{(t)})=\dim(\Delta_{{\mathbf{S}}\Theta}^{(t)})$ and $\dim(\xi_{\Lambda}^{(t)})=\dim(\xi_{{\mathbf{S}}\Theta}^{(t)})$, i.e., the output and states of our optimizers have the exact same dimension in both cases. This is not a coincidence: we now show that optimizing this linear adapter when $\Lambda$ is initialized to ${\mathbf{0}}$ is equivalent to optimizing $\Theta$ in the original neural network with linearly transformed gradient dynamics. (All proofs are in App. A).
@@
The above theorem holds for any optimizer of the form in eq. 2, e.g, Adam (Kingma & Ba, 2015). Notably, AdamW (Loshchilov & Hutter, 2019) does not fit this definition due to the way that weight decay is applied. See App. B for a discussion about weight decay, and what adjustments are required to preserve the equivalence.
@@ 3.2 Kronecker-factored Gradient Transformations
The formulation in § 3.1 assumes very little about the neural network being trained and the gradient transformation (or, equivalently, linear adapter) being applied, which makes it difficult to enable practical memory savings. Concretely, consider applying an arbitrary linear transformation to just a single linear layer of a neural network with parameters ${\mathbf{W}}\in\mathbb{R}^{m\times n}$, i.e., $f({\mathbf{X}};\Theta)={\mathbf{W}}{\mathbf{X}}$. In this case we have $\Theta=\operatorname{vec}\left({\mathbf{W}}\right)\in\mathbb{R}^{mn}$, and thus arbitrary linear maps of the form ${\mathbf{S}}\in\mathbb{R}^{r\times mn}$ require $O(mnr)$ memory to store. This cost is already non-trivial for a single linear layer of moderate size, and becomes rapidly intractable if we consider applying gradient transformations to the entirety of a model’s parameters. As such, practical applications need to consider matrices ${\mathbf{S}}$ that are efficient to store in memory (and also efficient to apply to $\Theta$).
To this end, we consider Kronecker-factored linear maps of the form ${\mathbf{S}}={\mathbf{R}}^{\top}\otimes{\mathbf{L}}$ where ${\mathbf{L}}\in\mathbb{R}^{d_{L}\times m},{\mathbf{R}}\in\mathbb{R}^{n\times d_{R}},d_{L}d_{R}=r$. This particular parameterization of ${\mathbf{S}}$ reduces the memory requirement to $O(d_{L}m+nd_{R})$ and FLOPs to $\min\{O(d_{L}mn+d_{L}nd_{R}),\,O(mnd_{R}+d_{L}md_{R})\}$ (since ${\mathbf{S}}\overline{\Theta}=\operatorname{vec}\left({\mathbf{L}}\overline{{\mathbf{W}}}{\mathbf{R}}\right)$), which can be memory-efficient if $d_{L},d_{R}$ are small enough. We now show applying § 3.1 to such an ${\mathbf{S}}$ establishes an equivalence between training with gradients transformed by ${{\mathbf{L}}\overline{{\mathbf{W}}}{\mathbf{R}}}$, and reparameterizing the linear layer as ${\mathbf{W}}+{\mathbf{L}}^{\top}{\mathbf{A}}{\mathbf{R}}^{\top}$ and only training ${\mathbf{A}}\in\mathbb{R}^{d_{L}\times d_{R}}$.
@@
§ 3.2 shows that MoRA (Jiang et al., 2024), LoRA-XS (Bałazy et al., 2024), and PMSS (Wang et al., 2025), which are recent approaches to parameter-efficient finetuning which reparameterize a linear layer as ${\mathbf{W}}+{\mathbf{B}}{\mathbf{A}}{\mathbf{C}}$ and only train $\mathbf{A}$, can be interpreted as training the model with linearly-transformed gradients where the linear transformation has a Kronecker factorization.
@@
Table 1: A summary of methods tested for our pretraining experiments, where we list the gradient transformation method (which is not relevant for Baseline/ReLoRA) and the corresponding adapter parameterization. We also break down the reparameterized model into trained and frozen components, alongside the the set of components that need to be persisted in memory (for methods that make use of easy-to-materialize random sketching matrices, viz., Gaussian, Rademacher, one only needs to persist the seeds for the gradient transformation). Random semi-orthogonal matrices—a tall/wide matrix whose columns/rows are orthonormal vectors—are also random but are not straightforwardly materializable from a seed, and hence may need to be persisted across optimization steps. In the Gaussian and Rademacher cases, we use $k$ as shorthand for the constant that rescales the distribution so that $\operatorname{\mathbb{E}}[{\mathbf{P}}{\mathbf{P}}^{\top}]={\mathbf{I}}$.
@@
[GaLore is one-sided LoRA]corollarygaloreIsLora Let ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ be the parameter matrix of a linear layer with corresponding gradient matrix $\overline{{\mathbf{W}}}\in\mathbb{R}^{m\times n}$. Without loss of generality, assume $m\leq n$. Now consider training ${\mathbf{W}}$ with $\operatorname{Optimizer}$ using GaLore, i.e., where we linearly transform the gradient matrix with a matrix ${\mathbf{P}}$ and then apply our optimizer on it, before transforming our update back to parameter space via ${\mathbf{P}}^{\top}$, viz.,
@@
The original GaLore work advocates for swapping out the gradient transformation every 200 optimizer steps. This does not break the equivalence in § 3.2. In the adapter formulation, recomputing the gradient transformation corresponds to merging the learned adapter into the frozen weights, updating the frozen part of the adapter, and resetting the learned part to zero. This effectively amounts to ReLoRA (Lialin et al., 2024), where one side of the adapter is kept frozen throughout training.
While §§ 3.2 and 3.2 focus on the case of a single linear layer, it is straightforward to generalize them to multiple linear layers. For example, one could treat the parameters of all layers as a single vector living in the product space of the individual layers’ parameter spaces, and define the gradient transformation map ${\mathbf{S}}$ on that space as applying the correct projection to each of the layer’s parameters individually. In practice, this can be implemented by modifying the optimizer step to apply a separate linear transformation to each layer.
Finally, we note that Hao et al. (2024) and Loeschcke et al. (2024) also show that training LoRA adapters with one side frozen with ordinary SGD is equivalent to applying a linear transformation to the gradient matrix, as in § 3.2. Our § 3.1 can be thus be seen as a generalization of these recent results, where we show that this equivalence generalizes to arbitrary parameters of the neural network and other types of stateful optimizers.

### v2

@@ Abstract
We study memory-efficient optimization of neural networks (in particular language models) with linear gradient transformations, where the gradients are linearly mapped to a lower dimensional space than the full parameter space, thus saving memory required for gradient accumulation and optimizer state persistence. The model parameters are updated by first performing an optimization step in the lower dimensional space and then going back into the original parameter space via the linear map’s transpose. We show that optimizing the model in this transformed space is equivalent to reparameterizing the original model through a linear adapter that additively modifies the model parameters, and then only optimizing the adapter’s parameters. When the transformation is Kronecker-factored, this establishes an equivalence between GaLore (Zhao et al., 2024) and one-sided LoRA (Hu et al., 2022). We show that this duality between gradient transformations and adapter-based reparameterizations unifies existing approaches to memory-efficient training and suggests new techniques for improving training efficiency and memory use.
@@ 1 Introduction
Training neural networks, in particular large language models (LLMs), can be extremely memory-intensive. Standard approaches for LLM training use gradient accumulation across multiple batches and optimizers such as Adam (Kingma & Ba, 2015), which maintains estimates of the first and second moments of the (stochastic) gradient. Hence, the amount of GPU memory needed for standard training can be as much as four times the amount of memory needed to store the model (assuming the gradients/optimization states are kept in the same precision as the model parameters).
@@
Our empirical experiments study this connection between linear gradient transformations and adapter-based reparameterizations in the context of memory-efficient LLM training. First, we perform a comparison across gradient projection-based and LoRA-based approaches for memory-efficient training and find that randomly sketching gradients works particularly well (§˜4.1). We also exploit the adapter view of projected-gradient training [^1] Note that this notion of performing gradient descent with projections of the gradient is distinct from projected gradient descent (PGD) from the optimization literature. by developing a QLoRA-style (Dettmers et al., 2023) approach to GaLore-style training. Second, we show that the gradient projection view of LoRA adapters can improve distributed training of LLMs with parallel LoRA adapters (Huh et al., 2024) by suggesting an initialization scheme of worker-specific LoRA adapters tailored for distributed training (§˜4.2). These results collectively demonstrate that this duality between linear gradient transformations and adapter-based reparameterizations is a productive lens with which to view neural network optimization, since it unifies several existing approaches and suggests new techniques for improving training efficiency and performance.
@@ 2.1 Memory Characteristics of LLM Training
Model parameters. We must keep the model’s parameters in memory, since these are used in various stages of the training process (e.g., to compute gradients). Here, it is useful to distinguish trainable parameters (which get updated regularly during training) from non-trainable parameters (which are not updated during training but may still be used in gradient computation).
@@
This work is mostly concerned with training LLMs in memory-constrained regimes where the model, optimizer, and gradient memory dominate, since activation storage can be made small by, e.g., reducing the microbatch size. As such, we will only focus on those categories in our calculations.
@@ 2.2 LoRA and GaLore
GaLore provides an alternative approach to memory-efficient pretraining. Instead of reparameterizing the weights to be a combination of full-rank and low-rank matrices—which increases the number of model parameters—GaLore performs a low-rank compression of the gradient matrix $\overline{{\mathbf{W}}}$ instead. The optimizer update is computed in this lower dimensional space, transformed back to the original space, and only then applied to the parameters. Specifically, given a gradient matrix $\overline{{\mathbf{W}}}\in\mathbb{R}^{m\times n}$, GaLore uses a matrix ${\mathbf{P}}\in\mathbb{R}^{k\times m}$ (with $k<m$) to transform the gradient via ${\mathbf{P}}\overline{{\mathbf{W}}}\in\mathbb{R}^{k\times n}$, feeds this compressed gradient into a regular optimizer to obtain a pseudo-parameter update $\Delta\in\mathbb{R}^{k\times n}$, and then updates the original parameters via ${\mathbf{P}}^{\top}{\Delta}$. In practice, $\mathbf{P}$} is given by the top singular vectors of $\overline{{\mathbf{W}}}$, where in order to amortize the cost of SVD, $\mathbf{P}$} is updated only every so often. As with LoRA, GaLore reduces the memory needed to store the optimizer states, since optimization happens in the lower dimensional space.
@@ 3 Duality between Linear Gradient Transformations and Adapters
In this section, we prove that training a neural network using linear transformations of the gradient is equivalent to reparameterizing the neural network using specific linear adapters. We begin with the general case, where all parameters are treated as arbitrary vectors (§˜3.1). We then show how applying a Kronecker-factored linear transformation to the gradients of linear layers of the network is equivalent to training the model with a version of LoRA which inserts a trainable matrix between the LoRA matrices (§˜3.2). From this, we further show that specializing to a specific choice of Kronecker-factored transformation establishes an equivalence between GaLore (Zhao et al., 2024) and “one-sided” LoRA (Hu et al., 2022) where one of the LoRA matrices is initialized in a particular way and kept frozen, while only the other one is trained (§˜3.2); this recovers the equivalence established in recent work (Hao et al., 2024; Loeschcke et al., 2024).
@@ 3.1 General Case
Let $f({\mathbf{X}};\Theta)$ be a neural network over input $\mathbf{X}$} with trainable parameters $\Theta\in\mathbb{R}^{d}$, and further let $\overline{\Theta}\in\mathbb{R}^{d}$ be the gradient of some differentiable loss function $\mathcal{L}$} of the network $f$ with respect to $\Theta$, computed on a random data minibatch. We use the superscript $\left(\cdot\right)^{(t)}$ to specify a particular quantity’s value after $t$ optimizer steps, e.g., $\Theta^{(t)}$ are the network’s parameters after $t$ optimizer steps. We are also interested in the optimization trajectory of a model $(\Theta^{(0)},\Theta^{(1)},\ldots,\Theta^{(t)})$. Our results show that two optimization trajectories—one from training with linear gradient transformations, and one from training with linear adapters—are equivalent.
@@
Because the above is adapting a neural network indirectly via another vector $\Lambda$ that is linearly mapped to the original parameter space, we refer to this as using a linear adapter, akin to the usage of “adapter” in the parameter-efficient finetuning literature (Houlsby et al., 2019; Hu et al., 2022). Since ${\mathbf{S}}^{\top}\in\mathbb{R}^{d\times r}$, we have $\dim(\Delta_{\Lambda}^{(t)})=\dim(\Delta_{{\mathbf{S}}\Theta}^{(t)})$ and $\dim(\xi_{\Lambda}^{(t)})=\dim(\xi_{{\mathbf{S}}\Theta}^{(t)})$, i.e., the output and states of our optimizers have the exact same dimension in both cases. This is not a coincidence: we now show that optimizing this linear adapter when $\Lambda$ is initialized to $\mathbf{0}$} is equivalent to optimizing $\Theta$ in the original neural network with linearly transformed gradient dynamics. (For this and all subsequent proofs, refer to App.˜A).
@@
The above theorem holds for any optimizer of the form in eq.˜2, e.g, Adam (Kingma & Ba, 2015). Notably, AdamW (Loshchilov & Hutter, 2019) does not fit this definition due to the way that weight decay is applied. See App.˜B for a discussion about weight decay, and what adjustments are required to preserve the equivalence.
@@ 3.2 Kronecker-factored Gradient Transformations
The formulation in §˜3.1 assumes very little about the neural network being trained and the gradient transformation (or, equivalently, linear adapter) being applied, which makes it difficult to enable practical memory savings. Concretely, consider applying an arbitrary linear transformation to just a single linear layer of a neural network with parameters ${\mathbf{W}}\in\mathbb{R}^{m\times n}$, i.e., $f({\mathbf{X}};\Theta)={\mathbf{W}}{\mathbf{X}}$. In this case we have $\Theta=\operatorname{vec}\left({\mathbf{W}}\right)\in\mathbb{R}^{mn}$, [^5] Intuitively, $\operatorname{vec}\left(\cdot\right)$ sends a matrix to its vectorized form (i.e., stacks its columns into a vector), and $\operatorname{vec}^{-1}\left(\cdot\right)$ is its inverse (i.e., unstacks vector back into matrix form). and thus arbitrary linear maps of the form ${\mathbf{S}}\in\mathbb{R}^{r\times mn}$ require $O(mnr)$ memory to store. This cost is already non-trivial for a single linear layer of moderate size, and becomes rapidly intractable if we consider applying gradient transformations to the entirety of a model’s parameters. As such, practical applications need to consider matrices $\mathbf{S}$} that are efficient to store in memory (and also efficient to apply to $\Theta$).
To this end, we consider Kronecker-factored linear maps of the form ${\mathbf{S}}={\mathbf{R}}^{\top}\otimes{\mathbf{L}}$ where ${\mathbf{L}}\in\mathbb{R}^{d_{L}\times m},{\mathbf{R}}\in\mathbb{R}^{n\times d_{R}},d_{L}d_{R}=r$. This particular parameterization of $\mathbf{S}$} reduces the memory requirement to $O(d_{L}m+nd_{R})$ and FLOPs to $\min\{O(d_{L}mn+d_{L}nd_{R}),\,O(mnd_{R}+d_{L}md_{R})\}$ (since ${\mathbf{S}}\overline{\Theta}=\operatorname{vec}\left({\mathbf{L}}\overline{{\mathbf{W}}}{\mathbf{R}}\right)$), which can be memory-efficient if $d_{L},d_{R}$ are small enough. We now show applying §˜3.1 to such an $\mathbf{S}$} establishes an equivalence between training with gradients transformed by ${\mathbf{L}}\overline{{\mathbf{W}}}{\mathbf{R}}$}, and reparameterizing the linear layer as ${\mathbf{W}}+{\mathbf{L}}^{\top}{\mathbf{A}}{\mathbf{R}}^{\top}$ and only training ${\mathbf{A}}\in\mathbb{R}^{d_{L}\times d_{R}}$.
@@
§˜3.2 shows that MoRA (Jiang et al., 2024), LoRA-XS (Bałazy et al., 2024), and PMSS (Wang et al., 2025), which are recent approaches to parameter-efficient finetuning which reparameterize a linear layer as ${\mathbf{W}}+{\mathbf{B}}{\mathbf{A}}{\mathbf{C}}$ and only train $\mathbf{A}$, can be interpreted as training the model with linearly-transformed gradients where the linear transformation has a Kronecker factorization.
@@
Table 1: A summary of methods tested for our pretraining experiments, where we list the gradient transformation method (which is not relevant for Baseline/ReLoRA) and the corresponding adapter parameterization. We also break down the reparameterized model into trained and frozen components, alongside the the set of components that need to be persisted in memory; for methods that make use of easy-to-materialize random sketching matrices (e.g., Gaussian) one only needs to persist the random number generator seeds for the gradient transformation, saving memory. Random semi-orthogonal matrices—a tall/wide matrix whose columns/rows are orthonormal vectors—are also random but are not straightforwardly materializable from a seed, and hence may need to be persisted across optimization steps. In the Gaussian and Rademacher cases, we use $k$ as shorthand for the constant that ensures that $\operatorname{\mathbb{E}}[{\mathbf{P}}{\mathbf{P}}^{\top}]={\mathbf{I}}$.
@@
[GaLore is one-sided LoRA]corollarygaloreIsLora Let ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ be the parameter matrix of a linear layer with corresponding gradient matrix $\overline{{\mathbf{W}}}\in\mathbb{R}^{m\times n}$. Without loss of generality, assume $m\leq n$. Now consider training $\mathbf{W}$} with $\operatorname{Optimizer}$ using GaLore, i.e., where we linearly transform the gradient matrix with a matrix $\mathbf{P}$}, apply the optimizer, and transform the update via $\mathbf{P}$$^{\top}$, viz.,
@@
The original GaLore work advocates for swapping out the gradient transformation every 200 optimizer steps. This does not break the equivalence in §˜3.2. In the adapter formulation, recomputing the gradient transformation corresponds to merging the learned adapter into the frozen weights, updating the frozen part of the adapter, and resetting the learned part to zero. This effectively amounts to ReLoRA (Lialin et al., 2024), where one side of the adapter is kept frozen throughout training.
While §§˜3.2 and 3.2 focus on the case of a single linear layer, it is straightforward to generalize them to multiple linear layers. For example, one could treat the parameters of all layers as a single vector living in the product space of the individual layers’ parameter spaces, and define the gradient transformation map $\mathbf{S}$} on that space as applying the correct projection to each of the layers’ parameters individually. This can be implemented by modifying the optimizer step to apply a separate linear transformation to each layer.
Finally, we note that Hao et al. (2024) and Loeschcke et al. (2024) also show that training LoRA adapters with one side frozen with ordinary SGD is equivalent to applying a linear transformation to the gradient matrix, and Muhamed et al. (2024) further shows this for more general optimizers as in §˜3.2. Our §˜3.1 can be thus be seen as a generalization of these recent results, where we show that this equivalence generalizes to arbitrary parameters of the neural network.